OpenStudy (anonymous):

For the real-valued functions g(x)=x+2/x-5 and h(x)=2x+7, find the composition G o H?

OpenStudy (anonymous):

and specify its domain using interval notation

OpenStudy (mathteacher1729):

What part of the problem are you getting stuck on?

OpenStudy (anonymous):

g(h(x))=g(2x+7) =(2x+7)+(2/((2x+7)-5))

OpenStudy (anonymous):

so i got in the end

OpenStudy (anonymous):

(4x^2+8x-19)/(2x+7)

OpenStudy (anonymous):

domain is all x except x= -7/2 is this all right and how do i turn this into interval notation

OpenStudy (mathteacher1729):

For the domain you do NOT want 1. zero in the denominator 2. natural log of a negative number or zero 3. sqrt of a complex number Look at g(h(x)) or to say it another way " g of h of x" or to say it another way g 0 h Do you see any of these things( 1, 2 or 3 ) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so i got the right answer?

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

its (2x + 9)/(2x + 2)

OpenStudy (mathteacher1729):

g(x)=x+2/x-5 and h(x)=2x+7 So \[\huge g(h(x)) = \frac{(2x+7) + 2}{(2x+7)-5}\] which is equal to \[\huge g(h(x)) = \frac{2x+9}{2x+2}\]

OpenStudy (anonymous):

yes

OpenStudy (mathteacher1729):

Ok, it is taking me too long to type in the LaTeX. ;-p

OpenStudy (mathteacher1729):

The denominator cannot equal zero. So when is 2x + 2 = 0 ?

OpenStudy (anonymous):

its domain is all x such that 2x + 2 ≠ 0 --> x ≠ - 1

OpenStudy (mathteacher1729):

Yes. In interval notation you would write something like... \[\huge (-\infty, -1)\cup (-1, \infty)\] But it's much easier to just write \[x \neq 0\] :-p

OpenStudy (anonymous):

kk thanks

OpenStudy (mathteacher1729):

No prob.